# Mingle: Mixtures of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

### v1

@@ 4.1 Experimental Setup
Evaluation Metrics. We evaluate performance using average accuracy (ACC) and backward transfer (BWT) [36]. ACC measures the mean accuracy over all tasks using the final merged model: $\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{t}^{\text{merged}}),$ where $a_{i}(\cdot)$ denotes the accuracy on task $i$. BWT quantifies the extent of forgetting by comparing performance on earlier tasks before and after merging: $\text{BWT}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left[a_{i}(\theta_{t}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\right].$
Table 1: Comparative results of continual merging methods, reporting average accuracy (ACC) and backward transfer (BWT) over ten task orders (mean±$\pm$±std). Best results are in bold; second-best are underlined. Mingle denotes a lightweight variant.
Method ViT-B/32 ViT-B/16 ViT-L/14
8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pre-Trained 48.1 56.9 55.6 55.4 62.0 59.8 64.9 69.1 65.6
Fine-Tuned 90.4 89.3 89.8 92.4 91.3 91.6 94.3 93.4 93.5
C. Fine-Tuned 79.8 67.4 62.6 82.9 72.2 68.2 90.0 70.9 77.7
ACC (%) ↑
Average (SWA) [25] 66.3 $\pm$0.0 65.4 $\pm$0.0 61.1 $\pm$0.0 72.3 $\pm$0.0 69.7 $\pm$0.0 64.8 $\pm$0.0 80.0 $\pm$0.0 77.5 $\pm$0.0 71.1 $\pm$0.0
C. Task Arithmetic [23] 67.5 $\pm$0.0 66.5 $\pm$0.0 60.0 $\pm$0.0 77.1 $\pm$0.0 70.9 $\pm$0.6 64.2 $\pm$0.0 82.1 $\pm$0.0 77.9 $\pm$0.0 70.3 $\pm$0.0
C. Ties-Merging [82] 49.0 $\pm$10.2 66.2 $\pm$0.6 59.9 $\pm$0.7 66.8 $\pm$3.7 70.5 $\pm$0.8 63.0 $\pm$1.6 64.3 $\pm$7.0 78.0 $\pm$0.6 68.3 $\pm$0.9
MagMax-Ind [39] 70.7 $\pm$0.0 67.0 $\pm$0.0 61.2 $\pm$0.0 76.7 $\pm$1.8 67.0 $\pm$0.0 62.5 $\pm$0.0 83.4 $\pm$0.0 71.2 $\pm$0.0 71.2 $\pm$0.0
OPCM [68] 75.5 $\pm$0.5 71.9 $\pm$0.3 65.7 $\pm$0.2 81.8 $\pm$0.3 77.1 $\pm$0.5 70.3 $\pm$0.2 87.0 $\pm$0.4 83.5 $\pm$0.2 76.0 $\pm$0.2
Mingle 85.8 $\pm$0.8 81.6 $\pm$1.4 77.1 $\pm$2.0 88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 - - -
Mingle 85.0 $\pm$0.5 81.7 $\pm$1.0 77.1 $\pm$1.3 87.0 $\pm$0.6 84.7 $\pm$1.0 81.6 $\pm$1.3 91.4 $\pm$0.3 89.2 $\pm$0.1 83.6 $\pm$0.6
BWT (%) ↑
Average (SWA) [25] -11.5 $\pm$2.2 -8.0 $\pm$1.3 -7.1 $\pm$2.1 -9.7 $\pm$1.5 -7.1 $\pm$1.4 -7.3 $\pm$1.7 -7.3 $\pm$1.4 -5.8 $\pm$1.0 -6.4 $\pm$1.5
C. Task Arithmetic [23] -9.6 $\pm$1.5 -1.3 $\pm$1.6 -3.4 $\pm$1.0 -4.2 $\pm$1.0 -1.3 $\pm$0.4 -3.6 $\pm$0.4 -7.1 $\pm$0.8 -1.8 $\pm$0.3 -3.3 $\pm$0.3
C. Ties-Merging [82] -15.3 $\pm$8.0 1.9 $\pm$0.6 -1.5 $\pm$0.7 -5.5 $\pm$0.4 1.4 $\pm$0.7 -1.5 $\pm$1.2 -13.0 $\pm$5.7 -1.1 $\pm$0.4 -2.9 $\pm$1.0
MagMax-Ind [39] -8.3 $\pm$1.3 -7.4 $\pm$1.4 -7.2 $\pm$1.6 -6.1 $\pm$1.3 -7.4 $\pm$2.0 -8.0 $\pm$2.2 -5.0 $\pm$0.8 -6.0 $\pm$2.1 -6.5 $\pm$2.1
OPCM [68] -6.3 $\pm$1.1 -6.0 $\pm$1.0 -7.8 $\pm$1.5 -4.8 $\pm$0.7 -5.1 $\pm$1.4 -6.3 $\pm$2.2 -2.6 $\pm$1.0 -4.3 $\pm$0.7 -6.5 $\pm$1.8
Mingle -0.6 $\pm$0.4 -1.1 $\pm$0.3 -2.2 $\pm$0.8 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4 - - -
Mingle -0.1 $\pm$0.1 -0.4 $\pm$0.1 -1.3 $\pm$0.6 -0.1 $\pm$0.1 -0.3 $\pm$0.1 -1.0 $\pm$0.4 -0.2 $\pm$0.0 -0.4 $\pm$0.2 -1.5 $\pm$0.6
Table 2: Robustness results for ViT-B/32 continual merged on four tasks.
ACC(%)
C. Layer-wise AdaMerging [83] 56.0 $\pm$5.3 47.5 $\pm$4.4 43.1 $\pm$2.3 43.3 $\pm$3.4 18.1 $\pm$4.7 46.6 $\pm$3.0 48.9 $\pm$4.8 49.1 $\pm$4.0 44.9
Mingle 89.9 $\pm$0.4 82.8 $\pm$0.8 67.5 $\pm$2.0 70.7 $\pm$1.2 37.9 $\pm$0.4 77.0 $\pm$0.7 80.1 $\pm$0.8 82.9 $\pm$0.9 73.2
BWT(%)
C. Layer-wise AdaMerging [83] -38.0 $\pm$7.1 -37.3 $\pm$5.9 -22.2 $\pm$3.0 -25.2 $\pm$4.5 -20.8 $\pm$6.3 -28.6 $\pm$4.0 -34.7 $\pm$6.5 -36.1 $\pm$5.3 -29.5
Mingle -0.2 $\pm$0.2 -0.1 $\pm$0.4 0.7 $\pm$1.0 0.6 $\pm$1.1 -0.2 $\pm$1.1 0.2 $\pm$0.7 0.0 $\pm$0.5 0.5 $\pm$0.5 -0.2
@@ 4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms existing continual merging approaches across all CLIP backbones and task counts, achieving the highest accuracy while nearly eliminating backward forgetting. These results confirm that our method consistently delivers both strong task accuracy and long-term stability under diverse continual scenarios.
@@
Conventional CL (Sequential fine-tuned)
@@
Continual Merging (Independent fine-tuned)
@@
Robustness to Test-Time Distribution Shifts. Following prior work [83, 66], we evaluate Mingle on seven corruptions (motion blur, impulse noise, Gaussian noise, pixelate, spatter, contrast, JPEG) and report results in Tab. 2. It preserves high accuracy and near-zero or even positive BWT, outperforming all baselines, whereas direct application of SOTA TTA-based merging (WEMOE, AdaMerging) in a continual setting fails without tailored designs to continual setup.
Figure 3: Efficiency analysis on TTA time, trainable/full parameter count, and accuracy for CLIP-ViT-B/32 merged on eight tasks.
Configuration Time T. Param F. Param ACC
attn & mlp ($r=64$) 78 s 83.0 K 173.1 M 85.8
@@
Comparison with Conventional CL. Tab. 3 evaluates two CL paradigms on the MTIL benchmark: conventional CL, where each task model is fine-tuned from its immediate predecessor; and continual merging, which fine-tunes each task model independently before fusion, eliminating inter-model dependencies and enabling flexible task ordering and model reuse. Within the merging family, Mingle sets a new state-of-the-art, and when integrated into a sequential fine-tuning pipeline, it matches the performance of SOTA CL methods. This demonstrates both its strength as a fusion strategy and its versatility across different training regimes.
4.3 Ablation Results and Analysis
Ablation Study.
We explore the contribution of each component in Tab. 4. Row 1 shows a fixed-weight merging of low-rank experts as our baseline. In Row 2, adding TTA boosts ACC substantially but at the cost of worsening BWT with more tasks. Row 3 demonstrates that freezing earlier gates curbs forgetting while retaining ACC gains. Row 4 then applies null-space constraints, yielding further BWT improvements. Finally, Row 5 presents the full method with adaptive relaxation, which best harmonizes accuracy and long-term stability.
Computation and Parameter Efficiency.  Tab. 3 compares different expert insertion layers and rank settings for CLIP-ViT-B/32 on eight tasks. Embedding experts in both attn and mlp layers achieves the highest accuracy at the cost of increased TTA time and parameter overhead. A hybrid scheme (qkv & fc1, $r=64$) delivers a favorable balance. In the rank sweep, moving from $r=32$ to $64$ yields a clear ACC gain, while further increasing $r$ leads to diminishing or negative returns. Experiments were run on a single RTX 4090 GPU. These results highlight Mingle’s flexibility under diverse resource budgets.
Number of Seed Samples  The quantity of seed samples plays a central role in determining the reliability and efficiency of TTA. We examine how varying the number of seed samples per class impacts performance with CLIP ViT-B/16. As shown in Fig. 6, even a single sample per class yields strong results, with only minor gains from additional samples. This indicates that TTA is largely robust to sample quantity. We also observe reduced variance across task orders as more samples are used, motivating our choice of five samples per class as a balanced trade-off between performance and efficiency.
@@ 4.3 Ablation Results and Analysis
Figure 6: Analysis of seed sample quantity on performance.
Hyper-parameter of Gate.  We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4, 5(a)). As shown in Fig. 5(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task; performance saturates at $k=3$ (Fig. 5(c)), indicating that a small number of task-specific directions suffices.

### v2

@@ 4.1 Experimental Setup
Evaluation Metrics. We evaluate using average accuracy (ACC) and backward transfer (BWT) [37]. ACC is the mean accuracy of the final merged model across all tasks: $\text{ACC}=\tfrac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{t}^{\text{merged}})$, where $a_{i}(\cdot)$ is accuracy on task $i$. BWT measures forgetting by comparing performance on earlier tasks before vs. after the final merge: $\text{BWT}=\tfrac{1}{T-1}\sum_{i=1}^{T-1}\big[a_{i}(\theta_{T}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\big]$.
[…]
Method Assump. ViT-B/32 ViT-B/16 ViT-L/14
DM / DA 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pre-Trained –  /  – 48.1 56.9 55.6 55.4 62.0 59.8 64.9 69.1 65.6
Fine-Tuned –  /  – 90.4 89.3 89.8 92.4 91.3 91.6 94.3 93.4 93.5
C. Fine-Tuned –  /  – 79.8 67.4 62.6 82.9 72.2 68.2 90.0 70.9 77.7
ACC (%) $\uparrow$ Average (SWA) [26]  /  66.3 $\pm$0.0 65.4 $\pm$0.0 61.1 $\pm$0.0 72.3 $\pm$0.0 69.7 $\pm$0.0 64.8 $\pm$0.0 80.0 $\pm$0.0 77.5 $\pm$0.0 71.1 $\pm$0.0
C. Task Arithmetic [24]  /  67.5 $\pm$0.0 66.5 $\pm$0.0 60.0 $\pm$0.0 77.1 $\pm$0.0 70.9 $\pm$0.6 64.2 $\pm$0.0 82.1 $\pm$0.0 77.9 $\pm$0.0 70.3 $\pm$0.0
C. Ties-Merging [86]  /  49.0 $\pm$10.2 66.2 $\pm$0.6 59.9 $\pm$0.7 66.8 $\pm$3.7 70.5 $\pm$0.8 63.0 $\pm$1.6 64.3 $\pm$7.0 78.0 $\pm$0.6 68.3 $\pm$0.9
MagMax-Ind [41]  /  70.7 $\pm$0.0 67.0 $\pm$0.0 61.2 $\pm$0.0 76.7 $\pm$1.8 67.0 $\pm$0.0 62.5 $\pm$0.0 83.4 $\pm$0.0 71.2 $\pm$0.0 71.2 $\pm$0.0
Consensus TA [77]  /  67.1 $\pm$0.4 64.1 $\pm$0.8 45.8 $\pm$1.5 72.8 $\pm$0.5 69.0 $\pm$0.0 49.9 $\pm$1.9 80.4 $\pm$0.5 75.0 $\pm$1.0 51.3 $\pm$2.4
OPCM [70]  /  75.5 $\pm$0.5 71.9 $\pm$0.3 65.7 $\pm$0.2 81.8 $\pm$0.3 77.1 $\pm$0.5 70.3 $\pm$0.2 87.0 $\pm$0.4 83.5 $\pm$0.2 76.0 $\pm$0.2
C. LW AdaMerging [87]  /  53.4 $\pm$ 3.2 59.8 $\pm$ 1.6 59.7 $\pm$ 7.4 59.9 $\pm$2.3 64.3 $\pm$ 1.2 61.5 $\pm$ 1.1 68.8 $\pm$ 2.9 73.1 $\pm$ 5.7 66.9 $\pm$ 1.1
C. LoRA-WEMOE [68]  /  68.8 $\pm$ 7.8 63.8 $\pm$ 3.4 49.6 $\pm$ 15.4 72.6 $\pm$ 3.7 67.9 $\pm$ 2.9 55.0 $\pm$ 7.0 75.6 $\pm$ 7.8 74.0 $\pm$ 5.0 56.9 $\pm$ 19.8
Mingle (Ours)  /  85.8 $\pm$0.8 81.6 $\pm$1.4 77.1 $\pm$2.0 88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 91.8 $\pm$0.2 88.8 $\pm$0.7 85.5 $\pm$1.3
Mingle (Ours)  /  85.0 $\pm$0.5 81.7 $\pm$1.0 77.1 $\pm$1.3 87.0 $\pm$0.6 84.7 $\pm$1.0 81.6 $\pm$1.3 91.4 $\pm$0.3 89.2 $\pm$0.1 83.6 $\pm$0.6
BWT (%) $\uparrow$ Average (SWA) [26]  /  -11.5 $\pm$2.2 -8.0 $\pm$1.3 -7.1 $\pm$2.1 -9.7 $\pm$1.5 -7.1 $\pm$1.4 -7.3 $\pm$1.7 -7.3 $\pm$1.4 -5.8 $\pm$1.0 -6.4 $\pm$1.5
C. Task Arithmetic [24]  /  -9.6 $\pm$1.5 -1.3 $\pm$1.6 -3.4 $\pm$1.0 -4.2 $\pm$1.0 -1.3 $\pm$0.4 -3.6 $\pm$0.4 -7.1 $\pm$0.8 -1.8 $\pm$0.3 -3.3 $\pm$0.3
C. Ties-Merging [86]  /  -15.3 $\pm$8.0 1.9 $\pm$0.6 -1.5 $\pm$0.7 -5.5 $\pm$0.4 1.4 $\pm$0.7 -1.5 $\pm$1.2 -13.0 $\pm$5.7 -1.1 $\pm$0.4 -2.9 $\pm$1.0
MagMax-Ind [41]  /  -8.3 $\pm$1.3 -7.4 $\pm$1.4 -7.2 $\pm$1.6 -6.1 $\pm$1.3 -7.4 $\pm$2.0 -8.0 $\pm$2.2 -5.0 $\pm$0.8 -6.0 $\pm$2.1 -6.5 $\pm$2.1
Consensus TA [77]  /  3.8 $\pm$0.9 -1.3 $\pm$0.9 -11.8 $\pm$1.9 3.5 $\pm$0.6 -1.1 $\pm$0.8 -11.6 $\pm$1.3 2.4 $\pm$0.6 -2.5 $\pm$0.8 -16.5 $\pm$1.5
OPCM [70]  /  -6.3 $\pm$1.1 -6.0 $\pm$1.0 -7.8 $\pm$1.5 -4.8 $\pm$0.7 -5.1 $\pm$1.4 -6.3 $\pm$2.2 -2.6 $\pm$1.0 -4.3 $\pm$0.7 -6.5 $\pm$1.8
C. LW AdaMerging [87]  /  -32.5 $\pm$3.6 -24.1 $\pm$1.7 -22.7 $\pm$4.3 -27.8 $\pm$2.7 -22.1 $\pm$1.4 -21.4 $\pm$1.2 -24.3 $\pm$3.3 -19.6 $\pm$1.7 -21.7 $\pm$1.1
C. LoRA-WEMOE [68]  /  -20.4 $\pm$9.0 -20.2 $\pm$3.9 -24.5 $\pm$10.0 -18.0 $\pm$6.2 -18.8 $\pm$3.4 -25.8 $\pm$7.9 -17.8 $\pm$5.9 -16.8 $\pm$5.3 -27.9 $\pm$17.2
Mingle (Ours)  /  -0.6 $\pm$0.4 -1.1 $\pm$0.3 -2.2 $\pm$0.8 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4 -0.6 $\pm$0.1 -1.0 $\pm$0.3 -2.6 $\pm$0.9
Mingle (Ours)  /  -0.1 $\pm$0.1 -0.4 $\pm$0.1 -1.3 $\pm$0.6 -0.1 $\pm$0.1 -0.3 $\pm$0.1 -1.0 $\pm$0.4 -0.2 $\pm$0.0 -0.4 $\pm$0.2 -1.5 $\pm$0.6
Table 2: Results of continual merging Flan-T5-base models on 8 tasks, ordered alphabetically.
Method DM / DA CoLA MNLI MRPC QNLI QQP RTE SST2 STSB ACC $\uparrow$ BWT $\uparrow$
Pre-trained –  /  – 69.1 56.5 76.2 88.4 82.1 80.1 91.2 62.2 75.7 -
Individual –  /  – 75.0 83.4 87.5 91.5 85.4 85.9 93.6 88.7 86.4 -
Task Arithmetic  /  69.1 58.1 77.9 88.9 83.1 79.1 90.7 74.0 77.6 -4.6
Ties-Merging  /  39.3 70.0 82.4 88.8 81.8 75.8 89.7 76.8 75.6 -6.1
OPCM  /  69.9 72.9 78.7 90.3 83.8 83.0 92.2 73.7 80.6 -2.5
LW AdaMerging  /  69.1 58.1 77.9 88.9 83.1 79.1 90.7 74.2 77.6 -4.7
LoRA-WEMOE  /  71.5 80.6 78.2 90.3 82.7 80.5 91.3 76.2 81.4 0.1
Mingle (Ours)  /  75.0 78.2 86.0 90.9 84.2 80.5 92.5 78.8 83.3 0.1
@@ 4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms prior continual merging methods on all CLIP backbones and task counts. It achieves the highest accuracy with backward forgetting kept near zero, demonstrating both strong forward learning and long-term stability. The lightweight variant performs on par with the full version, further underscoring the robustness of our approach. On NLP benchmarks (Tab. 2), Mingle likewise attains the best overall accuracy and non-negative BWT, improving on multiple GLUE tasks while maintaining balanced performance across the suite. Together, these results across vision and language confirm that Mingle consistently delivers state-of-the-art accuracy while nearly eliminating forgetting under diverse continual scenarios.
@@
Conventional CL (Sequential fine‐tuned)
@@
Continual Merging (Independent fine‐tuned)
@@
Table 4: Robustness results of ViT-B/32 continually merged across 4 tasks.
ACC (%) $\uparrow$ C. LW AdaMerging [87] 56.0 $\pm$5.3 47.5 $\pm$4.4 43.1 $\pm$2.3 43.3 $\pm$3.4 18.1 $\pm$4.7 46.6 $\pm$3.0 48.9 $\pm$4.8 49.1 $\pm$4.0 44.9
C. LoRA-WEMOE [68] 78.7 $\pm$4.5 71.0 $\pm$4.9 55.0 $\pm$3.8 59.4 $\pm$3.8 24.9 $\pm$24.9 60.5 $\pm$3.8 68.5 $\pm$4.8 69.7 $\pm$4.4 61.0
Mingle (Ours) 89.9 $\pm$0.4 82.8 $\pm$0.8 67.5 $\pm$2.0 70.7 $\pm$1.2 37.9 $\pm$0.4 77.0 $\pm$0.7 80.1 $\pm$0.8 82.9 $\pm$0.9 73.2
BWT (%) $\uparrow$ C. LW AdaMerging [87] -38.0 $\pm$7.1 -37.3 $\pm$5.9 -22.2 $\pm$3.0 -25.2 $\pm$4.5 -20.8 $\pm$6.3 -28.6 $\pm$4.0 -34.7 $\pm$6.5 -36.1 $\pm$5.3 -29.5
C. LoRA-WEMOE [68] -13.6 $\pm$6.9 -14.6 $\pm$8.2 -11.3 $\pm$4.7 -10.2 $\pm$3.7 -15.6 $\pm$8.8 -10.8 $\pm$3.9 -11.2 $\pm$8.8 -16.7 $\pm$6.6 -13.0
Mingle (Ours) -0.2 $\pm$0.2 -0.1 $\pm$0.4 0.7 $\pm$1.0 0.6 $\pm$1.1 -0.2 $\pm$1.1 0.2 $\pm$0.7 0.0 $\pm$0.5 0.5 $\pm$0.5 -0.2
Comparison with Conventional CL. Tab. 3 evaluates two CL paradigms on the MTIL benchmark: conventional CL, where each task model is fine-tuned from its immediate predecessor; and continual merging, which fine-tunes each task model independently before fusion, eliminating inter-model dependencies and enabling flexible task ordering and model reuse. Within the merging family, Mingle sets a new state‑of‑the‑art, and when integrated into a sequential fine‑tuning pipeline, it matches the performance of SOTA CL methods. This demonstrates both its strength as a fusion strategy and its versatility across different training regimes.
Robustness to Test-Time Distribution Shifts. Following prior work [87, 68], we evaluate Mingle on seven corruptions (motion blur, impulse noise, Gaussian noise, pixelate, spatter, contrast, JPEG) and report results in Tab. 4. It preserves high accuracy and near‑zero or even positive BWT, outperforming all baselines, whereas direct application of SOTA TTA‑based merging (WEMOE, AdaMerging) in a continual setting fails without tailored designs to continual setup.
4.3 Ablation Results and Analysis
Ablation Study. We explore the contribution of each component in Tab. 5. Row 1 shows a fixed‑weight merging of low‑rank experts as our baseline. In Row 2, adding TTA boosts ACC substantially but at the cost of worsening BWT with more tasks. Row 3 demonstrates that freezing earlier gates curbs forgetting while retaining ACC gains. Row 4 then applies null‑space constraints, yielding further BWT improvements. Finally, Row 5 presents the full method with adaptive relaxation, which best harmonizes accuracy and long‑term stability.
@@ 4.3 Ablation Results and Analysis
Table 6: Ablation on number of adaptation steps for ViT-B/32 across 8, 14, and 20 tasks.
Steps ACC (8-task) BWT (8-task) ACC (14-task) BWT (14-task) ACC (20-task) BWT (20-task)
[…]
Table 7: Efficiency and sample analysis. (a) Expert insertion layers and rank sweep over 8 tasks on CLIP-ViT-B/32. (b) Wall-clock adaptation time across tasks on CLIP-ViT-B/32. (c) Accuracy (%) of CLIP-ViT-B/16 under varying numbers of test samples.
(a) Expert insertion layers and rank sweep (8 tasks)
Configuration TTA Time Train. Param Full Param ACC(%)
@@
qkv & fc1 ($r=64$) 65 s 36.9 k 139.7 M 85.0
(b) Wall-clock adaptation time across tasks
#Tasks Adaptation steps Total Time (s) Avg./task (s)
8 50 78 9.8
14 50 138 9.9
20 50 211 10.6
(c) Results with varying samples/class
# Samples/class 8-task 14-task 20-task
[…]
Ablation on TTA optimization steps Tab. 6 reports accuracy and backward transfer of CLIP ViT-B/32 under different adaptation steps across 8, 14, and 20 tasks. Accuracy improves steadily with longer schedules, rising from 60–63% at 5 steps to 77–86% at 50 steps. Forgetting remains negligible, with BWT close to zero in all cases and only a minor drop of about 2% in the 20-task setting at 50 steps. Notably, even 20 steps are sufficient to surpass all baselines in Tab. 1, while 50 steps provide the best accuracy at modest additional cost. These results confirm that Mingle is both effective under tight compute budgets and scalable with more adaptation.
Computation and Parameter Efficiency. Tab. 7 summarizes the efficiency analysis. In part (a), inserting experts into both attn and mlp layers yields the highest accuracy, 85.8%, but also the longest adaptation time of 78s and 83k trainable parameters. A lighter hybrid scheme qkv & fc1 with rank 64 reaches 85.0% accuracy with 36.9k parameters and 65s, offering a better trade-off. The rank sweep shows that raising the rank from 32 to 64 improves accuracy from 84.5% to 85.0%, while larger ranks bring little or even negative gain, e.g., rank 768 drops to 83.7%. Part (b) reports wall-clock adaptation time as tasks increase: with 20 tasks the total is 211s, averaging about 10s per task. After adaptation the router remains fixed and inference is purely feedforward without TTA, enabling low-latency deployment across all tasks. Overall, Mingle achieves a strong balance of accuracy, parameter efficiency, and scalability under diverse resource budgets.
Number of Seed Samples The number of seed samples per class is crucial for TTA reliability and efficiency. As shown in Tab. 7 (c), using no samples reduces to the Static baseline, where LoRA modules are merged with fixed coefficients (0.3) without adaptation, yielding 70–79% accuracy. Introducing a single sample lifts accuracy to 81–88%, and adding more samples offers only minor gains. Variance across task orders decreases with more samples, making five samples per class a balanced trade-off between performance and efficiency.
@@
Visualization on Gate Activations. Fig. 3 shows that the null-space constraint suppresses gate responses on previously seen tasks, reducing forgetting, with smaller $\gamma$ giving stronger attenuation. We also observe that gate activations remain below 1.0 across tasks, even in the w/o Null-Space variant (blue curve), showing that under-activation is not solely due to the constraint. Instead, it reflects the complementary nature of experts: multiple LoRA modules capture overlapping but distinct subspaces, and softly combining them often yields better performance, especially in settings like TTCMM where task boundaries are fuzzy.
Hyper-parameter of Gate. We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4(a)). As shown in Fig. 4(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task; performance saturates at $k=3$ (Fig. 4(c)), indicating that a small number of task-specific directions suffices.